Solve the simultaneous equation
3x+5y=21
3x-y=3

Answer by MathTutor16(2)

3x+5y=21 ...1
3x-y=3 ...2
The easiest way to solve this problem would be to get rid of one of the variables. In both equations you have 3x. If you subtract the two equations, you get the following:
0x+6y=18
6y=18
From this, it follows that y=3. If you now substitute this value into EITHER equation, you will be able to find x. Plugging into the second equation:
3x-3=3
3x=6
x=2
Therefore, the answer is x=2,y=3. To check, you can plug in these values into either equation:
3x+5y=21
(3*2)+(5*3)=21
6+15=21
21=21
This proves that our answer is correct.